# The universal multiplicity function: counting halos and voids

### v1

@@ 4 The effective moving barrier for cosmic voids
To compute the moving barrier for halos and voids we measure it using numerical realizations of 3-dimensional Gaussian random fields, representing the Lagrangian space. The procedure consists in finding halos and voids, following their definition in Sec. 2, to measure the corresponding multiplicity function, and to derive the moving barrier generating it (see Appendix D for details). In this work, we use as Lagrangian space the initial conditions (ICs) of cosmological simulations, which satisfy the above description. In this section we focus on voids alone. This is because, in Lagrangian space, the halo treatment is analogous to the void case, but with a positive formation threshold. In addition, in the literature, multiplicity function models for voids are not as advanced as the ones for halos.
Simulations and void finder
In this work, we use the standard “Dark Energy and Massive Neutrino Universe” (DEMNUni) set of simulations (Carbone et al., 2016; Parimbelli et al., 2022) and new high-resolution runs with 64 times better mass resolution (HR-DEMNUni) (see e.g. Hernández-Molinero et al., 2023). These simulations have been produced using the tree particle mesh-smoothed particle hydrodynamics (TreePM-SPH) code Gadget-3 (Springel, 2005), and are characterized by a Planck 2013 (Planck Collaboration et al., 2014) baseline flat $\Lambda$CDM cosmology with different values of the total neutrino mass and the parameters characterizing the dark energy equation of state. Here we consider only the massless neutrino flat $\Lambda$CDM case and use two runs characterized by volumes of $(500\,h^{-1}{\rm Mpc)^{3}}$ and $(2\,h^{-1}{\rm Gpc})^{3}$, respectively, both with $2048^{3}$ DM particles and, when present, $2048^{3}$ neutrino particles. In particular, we consider the initial matter density field of the $\Lambda$CDM simulation, at $z=99$, for both the the DEMNUni and HR-DEMNUni simulations. This set of simulations fits well with the goal of this work, as their ICs are produced at high redshift with the Zeld’ovich approximation, ensuring that the realization of the field is Gaussian and fully described by the linear theory. Moreover, their combination of volume and resolution ensure a statistically relevant sample of voids from small to large scales (Kreisch et al., 2019; Schuster et al., 2019; Verza et al., 2019, 2022b, 2022a; Vielzeuf et al., 2023).
@@
Table 1: Void formation threshold explored. The first column, $\delta(z_{\rm IC})$, lists the actual density contrast in the simulation ICs at $z_{\rm IC}=99$; the second column, $\delta_{\rm v}$, is the corresponding linear density contrast linearly evolved to $z=0$; the last two columns, $\delta^{\rm E}_{\rm v}(z=1)$ and $\delta^{\rm E}_{\rm v}(z=0)$, list the corresponding Eulerian (non-linear) density contrast evolved at $z=1,0$, assuming the spherical symmetry.
@@
To identify voids according to the definition given in Sec. 2, we used the Pylians3222https://github.com/franciscovillaescusa/Pylians3 (Villaescusa-Navarro, 2018) implementation of the void finding algorithm of Banerjee & Dalal (2016), therefore no post-processing of the void catalog is needed. We explore several void formation thresholds, listed in Tab. 1. The first column lists their actual values in the simulation ICs at $z=99$; the second one the corresponding values linearly evolved at $z=0$; the last two columns, $\delta^{\rm E}_{\rm v}(z=1)$ and $\delta^{\rm E}_{\rm v}(z=0)$, list the corresponding Eulerian (non-linear) density contrast evolved down to $z=1$ and 0, assuming the spherical symmetry (SVdW; Pace et al., 2010, 2017). The exact mapping from the Lagrangian to the Eulerian density contrast may be slightly different from the spherical evolution. Nevertheless, for voids this is an accurate approximation (Massara & Sheth, 2018, see Sec. 5). These values are chosen to span a wide range of nonlinear values, from $\delta^{\rm E}_{\rm v}(z=0)\sim-0.8$ to $\delta^{\rm E}_{\rm v}(z=0)\sim-0.3$. Note that the void sample target of galaxy surveys at $z\lesssim 2$ have a DM density contrast in a range between $-0.4$ and $-0.3$ (Contarini et al., 2022).
Methods and results
@@
To find the effective moving barrier for voids, we fit the moving barrier against the one derived from the Lagrangian VSF (i.e. the one measured in the ICs density field of DEMNUni and HR-DEMNUni simulations), considering the general functional shape (SMT; Sheth & Tormen, 2002; Paranjape et al., 2012; Sheth et al., 2013)
$B(\sigma)=\alpha\left[1+(\beta/\sigma)^{\gamma}\right].$ (16)
We perform a Bayesian MCMC analysis of the theoretical multiplicity function corresponding to the above three-parameter moving barrier, to fit $\alpha$, $\beta$, and $\gamma$ as explained in Appendix D.
@@
Figure 2: Projected posterior distribution (orange violins) of the $\alpha$, $\beta$, $\gamma$ parameters of Eq. (16), measured against the Lagrangian VSF in the DEMNUni and HR-DEMNUni ICs, as a function of the formation threshold threshold $\delta_{\rm v}$ (vertical orange dashed lines); each panel shows the corresponding projection. Blue solid lines and dashed areas show the best-fit and 68% CL of the scaling relations described in the text.
@@
In the above equation, $\alpha$ modulates the global barrier amplitude, so we expect a dependence from the void formation threshold. The $\beta/\sigma$ quantity has a clear relation with the threshold density in terms of the standard deviation of matter fluctuations, $\sigma$, usually defined as $\nu=|\delta_{\rm v}|/\sigma$. Therefore, a dependence of $\beta$ on $|\delta_{\rm v}|$ is expected. Finally, the $\gamma$ parameter controls the shape of the moving barrier. Fig. 2 shows the projected posterior distribution (orange violins) for the parameters in Eq. (16) resulting from each explored void formation threshold, $\delta_{\rm v}$, whose reported value is the one linearly extrapolated at $z=0$. We find that the parameters $\alpha$ and $\beta$ are consistently described by a linear relation in terms of the linear void formation threshold. The $\gamma$ parameter can be consistently described as constant at any formation threshold. Implementing a linear fit with a MCMC, we obtain
@@
$\ \alpha=m_{\alpha}|\delta_{\rm v}|+q_{\alpha}:\\ \beta=m_{\beta}|\delta_{\rm v}|+q_{\beta}:\end{matrix}\quad\begin{matrix}m\\ 0.517\pm^{0.031}_{0.035}\\ 0.098\pm^{0.032}_{0.034}\end{matrix}\ \left|\ \begin{matrix}q\\ -0.089\pm^{0.031}_{0.036}\\ 0.103\pm^{0.053}_{0.043}\end{matrix}\right.\\ &\gamma=0.87\pm^{0.07}_{0.07}.\end{split}$ (17)
@@
Figure 3: Lagrangian VSF for the linear threshold $\delta_{\rm v}=-0.623$. Top panel: measurements from the DEMNUni ICs (black dots), HR-DEMNUni ICs (red dots) with the corresponding Poissonian uncertainty (error bars); VSF from the effective barrier approach presented in this work, Eq. (16), with the MCMC best-fit parameters, Eq. (17). The solid blue line shows the exact numerical solution, Eqs. (5)–(6), the blue dashed line the analytical approximation, Eq. (3); VSF from the standard excursion-set approach with a top-hat filter, i.e. $B=\delta_{\rm v}$ in Eqs. (5)–(6) (orange line); SVdW model (green line). Bottom panel: residuals of the VSF obtained via the effective barrier best-fit with respect to the VSF measured in the DEMNUni ICs (black lines) and HR-DEMNUni ICs (red lines), in units of the Poissonian uncertainty. Solid lines shows the exact numerical solution, dashed lines the corresponding analytical approximation. The shaded areas show the 68% CL; the hatched area shows ±$\pm 1\sigma$ interval.
@@
The upper panel of Fig. 3 shows the Lagrangian VSF with linear formation threshold $\delta_{\rm v}=-0.623$, as measured from the DEMNUni ICs (black dots) and HR-DEMNUni ICs (red dots), with the corresponding Poissonian uncertanty (errorbars), together with the following theoretical modelings: i) the VSF from the standard excursion-set multiplicity function, i.e. using a top-hat filter and a constant void formation threshold corresponding to the physical one, $B(S)=\delta_{\rm v}$ in Eqs. (5)–(6) (orange curve); ii) the Lagrangian VSF of SVdW (green curve) with the corresponding $\delta_{\rm v}=-0.623$: the global behavior is similar to the previous case, but its amplitude is slightly higher. This is because the SVdW is the exact first crossing double barrier solution with the sharp-$k$ filter, i.e. of Markovian random walks, for which the crossing probability is higher; iii) the VSF from the effective barrier approach presented in this work, Eq. (16), with the MCMC best-fit parameters, Eq. (17). The solid blue line shows the exact numerical solution, Eqs. (5)–(6), the blue dashed line the analytical approximation, Eq. (3). The shaded area shows the 68% CL interval of the posterior distribution of the parameters propagated to the VSF. Finally, the lower panel of Fig. 3 shows the residuals of the VSF obtained via our new multiplicity function and the effective barrier model with respect to VSF measurements from the DEMNUni ICs (black lines) and HR-DEMNUni ICS (red lines), in units of the Poissonian uncertainty. The solid lines shows the exact numerical solution, Eqs. (5)–(6), the dashed lines the analytical approximation, Eq. (3). The shaded areas show the 68% CL, while the hatched area shows ±$\pm 1\sigma$ interval. The agreement is within $1$-$\sigma$ on almost all the scales. The threshold $\delta_{\rm v}=-0.623$ is chosen as it is representative of the void population detectable around $z\sim 1$ in galaxy surveys (Contarini et al., 2022). The results of the other thresholds explored are qualitatively analogous. Eqs. (16)–(17) represent the second most important result of our work.
@@
As final consideration here we have modeled the void multiplicity function with a single moving barrier. This may seam in contrast to what written in Sec. 2, i.e. that void formation is characterized by two barriers: the formation and the collapsing ones. The double barrier was introduced to account for the void-in-cloud effect, which however occurs at scales smaller than the ones explored in this work (SVdW). Moreover, the use of a top-hat filter and the focus on minima in the density field reasonably push the void-in-cloud effect at scales even smaller than in SVdW [^3] This last condition describes the filtered density fields as Markovian random walks corresponding to a large variation of the field at relatively high $\sigma$ (low $R$). This enhances the void-in-cloud fraction with respect to the correlated step case.. We verified this by measuring in the DEMNUni-ICs the density profile of each void up to 100 $h^{-1}$pc. We have not found any void embedded in large overdensities that reaches neither the linear spherical collapsing threshold extrapolated at $z=0$, $\delta_{c}\simeq 1.686$, nor the turn-around threshold, $\delta_{\rm ta}\simeq 1.06$.
@@ 5 Eulerian mapping
In this Section, we introduce the Lagrangian to Eulerian mapping assuming the simplest possible approximation: the spherical map. With this assumption, each Lagrangian halo and void is considered evolving following the spherically symmetric gravitational collapse/expansion (SVdW; Pace et al., 2010, 2017). For both the halo and the void cases, the number of objects is assumed to be conserved from Lagrangian to Eulerian space. Considering the void case, due to mass conservation, the corresponding Eulerian radius is $R_{\rm E}=(1+\delta_{\rm v}^{\rm E})^{-1/3}R$, where the super- and sub-script E denotes Eulerian quantities. For halos, the spherical symmetric collapse implies that all fluid elements of a Lagrangian halo collapse in the Eulerian space, therefore $M_{\rm E}(R)=M_{\rm TH}(R)$ as defined in Sec. 2. Note that the multiplicity function is a differential quantity; therefore, to express it as a function of Eulerian quantities, we use the relation
$f(S)\,{\rm d}S=f(R)\,{\rm d}R=f(R_{\rm E})\,{\rm d}R_{\rm E}=f(M_{\rm E})\,{\rm d}M_{\rm E}.$ (18)
Figure 4: Top panel: VSF measurements from DEMNUni (black dots) and HR-DEMNUni (red dots) at $z=0$, for voids reaching $\delta^{\rm E}_{\rm v}=-0.409$ threshold, corresponding to the Eulerian spherical map of the linear threshold $\delta_{\rm v}=-0.623$. Errorbars shows the Poissonian uncertainty; Eulerian spherical mapping, Eq. (19), of the VSF with the effective barrier approach implemented in this work, Eqs. (16)–(17) and $\delta_{\rm v}=-0.623$, for the numerical solution (blue solid), Eqs. (5)–(6). The blue dashed lines show the corresponding analytical approximation, Eq. (3). Eulerian spherical mapping, Eq. (19), of the SVdW model (green line). Vdn model (pink line). Bottom panel: relative difference between the theoretical models appearing in the upper panel (same colors) with respect to the VSF measured in DEMNUni (dark color) and HR-DEMNUni (transparent). The errorbars show the corresponding Poissonian uncertainty.
The cosmic void case. From the above equation it follows that the Eulerian multiplicity function is $f_{\rm E}(R_{\rm E})=f[R(R_{\rm E})]\,{\rm d}R/{\rm d}R_{\rm E}=(1+\delta_{\rm v}^{\rm E})^{1/3}f[R(R_{\rm E})]$. The Eulerian void size function, therefore, is
$\frac{{\rm d}n(R_{\rm E})}{{\rm d}R_{\rm E}}=\frac{3}{4\pi R^{3}(R_{\rm E})}f_{\rm E}(R_{\rm E}).$ (19)

### v2

@@ 4 The effective moving barrier for cosmic voids
To compute the moving barrier for haloes and voids we measure it using numerical realisations of 3-dimensional Gaussian random fields, representing the Lagrangian space. The procedure consists in finding haloes and voids, following their definition in Sec. 2, to measure the corresponding multiplicity function, and to derive the moving barrier generating it (see Appendix E for details). In this work, we use as Lagrangian space the initial conditions (ICs) of cosmological simulations, which satisfy the above description. In this section we focus on voids alone. This is because, in Lagrangian space, the halo treatment is analogous to the void case, but with a positive formation threshold. In addition, in the literature, multiplicity function models for voids are not as advanced as the ones for haloes.
4.1 Simulations and void finder
In this work, we use the standard “Dark Energy and Massive Neutrino Universe” (DEMNUni) set of simulations [49, 50] and new high-resolution runs with 64 times better mass resolution (HR-DEMNUni) [see e.g. 51]. These simulations have been produced using the tree particle mesh-smoothed particle hydrodynamics (TreePM-SPH) code Gadget-3 [52], and are characterised by a Planck 2013 [53] baseline flat $\Lambda$CDM cosmology with different values of the total neutrino mass and the parameters characterising the dark energy equation of state. Both the standard and HR runs involve $2048^{3}$ DM particles and, when present, $2048^{3}$ neutrino particles, in a volume of $(2000\,h^{-1}{\rm Mpc})^{3}$ and $(500\,h^{-1}{\rm Mpc})^{3}$, respectively. Here we consider only the massless neutrino flat $\Lambda$CDM case.
@@ 4.1 Simulations and void finder
Table 1: Void formation threshold explored. The first column, $\delta(z_{\rm IC})$, lists the actual density contrast in the simulation ICs at $z_{\rm IC}=99$; the second column, $\delta_{\rm v}$, is the corresponding linear density contrast, linearly evolved to $z=0$; the last two columns, $\delta^{\rm E}_{\rm v}(z=1)$ and $\delta^{\rm E}_{\rm v}(z=0)$, list the corresponding Eulerian (non-linear) density contrast evolved at $z=1,0$, assuming the spherical symmetry.
In particular, we consider the initial matter density field of the $\Lambda$CDM simulation, at $z=99$, for both the DEMNUni and HR-DEMNUni simulations. This set of simulations fits well with the goal of this work, as their ICs are produced at high redshift with the Zeld’ovich approximation, ensuring that the realisation of the field is Gaussian and fully described by the linear theory. Moreover, their combination of volume and resolution ensure a statistically relevant sample of voids from small to large scales [54, 55, 27, 56, 57, 58].
@@
To identify voids according to the definition given in Sec. 2, we used the Pylians3222https://github.com/franciscovillaescusa/Pylians3 [59] implementation of the void finding algorithm of [60], therefore no post-processing of the void catalogue is needed. We explore several void formation thresholds, listed in Tab. 1. The first column lists their actual values in the simulation ICs at $z=99$; the second one the corresponding values linearly evolved at $z=0$; the last two columns, $\delta^{\rm E}_{\rm v}(z=1)$ and $\delta^{\rm E}_{\rm v}(z=0)$, list the corresponding Eulerian (non-linear) density contrast evolved down to $z=1$ and 0, assuming the spherical symmetry [SVdW; 61, 62]. The exact mapping from the Lagrangian to the Eulerian density contrast may be slightly different from the spherical evolution. Nevertheless, for voids this is an accurate approximation [30, see Sec. 5]. These values are chosen to span a wide range of nonlinear values, from $\delta^{\rm E}_{\rm v}(z=0)\sim-0.8$ to $\delta^{\rm E}_{\rm v}(z=0)\sim-0.3$. Note that the void sample target of galaxy surveys at $z\lesssim 2$ have a DM density contrast in a range between $-0.4$ and $-0.3$ [28].
4.2 Methods and results
@@
To find the effective moving barrier for voids, we fit the moving barrier against the multiplicity function derived from the Lagrangian VSF (i.e. measured in the ICs density field of DEMNUni and HR-DEMNUni simulations), considering the general functional shape [SMT; 17, 19, 21]
@@
Figure 2: Projected posterior distribution (orange violins) of the $\alpha$, $\beta$, $\gamma$ parameters of Eq. (4.1), measured against the Lagrangian VSF in the DEMNUni and HR-DEMNUni ICs, as a function of the formation threshold threshold $\delta_{\rm v}$ (vertical orange dashed lines); each panel shows the corresponding projection. Blue solid lines and dashed areas show the best-fit and 68% CL of the scaling relations described in the text.
$B(\sigma)=\alpha\left[1+(\beta/\sigma)^{\gamma}\right].$ (4.1)
@@
We perform a Bayesian MCMC analysis of the theoretical multiplicity function corresponding to the above three-parameter moving barrier, to fit $\alpha$, $\beta$, and $\gamma$ as explained in Appendix E. In the above equation, $\alpha$ modulates the global barrier amplitude, so we expect a dependence from the void formation threshold. The $\beta/\sigma$ quantity has a clear relation with the threshold density in terms of the standard deviation of matter fluctuations, $\sigma=\sqrt{S}$, usually defined as $\nu=|\delta_{\rm v}|/\sigma$. Therefore, a dependence of $\beta$ on $|\delta_{\rm v}|$ is expected. Finally, the $\gamma$ parameter controls the shape of the moving barrier. Fig. 2 shows the projected posterior distribution (orange violins) for the parameters in Eq. (4.1) resulting from each explored void formation threshold, $\delta_{\rm v}$, whose reported value is the one linearly extrapolated at $z=0$. We find that the parameters $\alpha$ and $\beta$ are consistently described by a linear relation in terms of the linear void formation threshold. The $\gamma$ parameter can be consistently described as constant at any formation threshold. Implementing a linear fit with a MCMC, we obtain
@@
$\ \alpha=m_{\alpha}|\delta_{\rm v}|+q_{\alpha}:\\ \beta=m_{\beta}|\delta_{\rm v}|+q_{\beta}:\end{matrix}\quad\begin{matrix}m\\ 0.517\pm^{0.031}_{0.035}\\ 0.098\pm^{0.032}_{0.034}\end{matrix}\ \left|\ \begin{matrix}q\\ -0.089\pm^{0.031}_{0.036}\\ 0.103\pm^{0.053}_{0.043}\end{matrix}\right.\\ &\gamma=0.87\pm^{0.07}_{0.07}.\end{split}$ (4.2)
@@
Figure 3: Lagrangian VSF for the linear threshold $\delta_{\rm v}=-0.623$. Top panel: measurements from the DEMNUni ICs (black dots), HR-DEMNUni ICs (red dots) with the corresponding Poissonian uncertainty (error bars); VSF from the effective barrier approach presented in this work, Eq. (4.1), with the MCMC best-fit parameters, Eq. (4.2). The solid blue line shows the exact numerical solution, the blue dashed line the analytical multiplicity function, Eq. (3.14), and the blue dotted line the corresponding approximation, Eq. (3.2); VSF from the standard excursion-set approach with a top-hat filter, i.e. $B=\delta_{\rm v}$ in Eqs. (3.4)–(3.5) (orange line); SVdW model (green line). Bottom panel: residuals of the VSF obtained via the effective barrier best-fit with respect to the VSF measured in the DEMNUni ICs (black lines) and HR-DEMNUni ICs (red lines), in units of the Poissonian uncertainty. Solid lines shows the exact numerical solution, dashed lines the analytical multiplicity function, Eq. (3.14), and dotted line the corresponding approximation, Eq. (3.2).
@@
The upper panel of Fig. 3 shows the Lagrangian VSF with linear formation threshold $\delta_{\rm v}=-0.623$, as measured from the DEMNUni ICs (black dots) and HR-DEMNUni ICs (red dots), with the corresponding Poissonian uncertanty (errorbars), together with the following theoretical modellings: i) the VSF from the standard excursion-set multiplicity function, i.e. using a top-hat filter and a constant void formation threshold corresponding to the physical one, $B(S)=\delta_{\rm v}$ in Eqs. (3.4)–(3.5) (orange curve); ii) Lagrangian VSF of SVdW (green curve) with the corresponding $\delta_{\rm v}=-0.623$: the global behaviour is similar to the previous case, but its amplitude is slightly higher. This is because the SVdW multiplicity function is the exact first crossing double barrier solution with the sharp-$k$ filter, i.e. of Markovian random walks, for which the crossing probability is higher; iii) the VSF from the effective barrier approach presented in this work, Eq. (4.1), with the MCMC best-fit parameters, Eq. (4.2). The solid blue line shows the exact numerical solution, Eqs. (3.4)–(3.5), the blue dashed line the analytical multiplicity function, Eq. (3.14), and the blue dotted line the corresponding approximation, Eq. (3.2). The shaded area shows the 68% CL interval of the posterior distribution of the parameters propagated to the VSF. Finally, the lower panel of Fig. 3 shows the residuals of the VSF obtained via our new multiplicity function and the effective barrier model with respect to VSF measurements from the DEMNUni ICs (black lines) and HR-DEMNUni ICS (red lines), in units of the Poissonian uncertainty. The solid lines shows the exact numerical solution, Eqs. (3.4)–(3.5), dashed lines the analytical multiplicity function, Eq. (3.14), and blue dotted line the corresponding approximation, Eq. (3.2). The shaded areas show the 68% CL, while the hatched area shows ±$\pm 1\sigma$ interval. The agreement is within $1$-$\sigma$ on almost all the scales. The threshold $\delta_{\rm v}=-0.623$ is chosen as it is representative of the void population detectable around $z\sim 1$ in galaxy surveys [28]. The results of the other thresholds explored are qualitatively analogous. Eqs. (4.1)–(4.2) represent the second most important result of our work.
@@
As final consideration here we have modelled the void multiplicity function with a single moving barrier. This may seam in contrast to what written in Sec. 2, i.e. that void formation is characterised by two barriers: the formation and the collapsing ones. The double barrier was introduced to account for the void-in-cloud effect, which however occurs at scales smaller than the ones explored in this work [SVdW]. Moreover, the use of a top-hat filter and the focus on minima in the density field reasonably push the void-in-cloud effect at scales even smaller than in SVdW [^3] This last condition describes the filtered density fields as Markovian random walks corresponding to a large variation of the field at relatively high $\sigma$ (low $R$). This enhances the void-in-cloud fraction with respect to the correlated step case.. We verified this by measuring in the DEMNUni-ICs the density profile of each void up to 100 $h^{-1}$pc. We have not found any void embedded in large overdensities that reaches neither the linear spherical collapsing threshold extrapolated at $z=0$, $\delta_{c}\simeq 1.686$, nor the turn-around threshold, $\delta_{\rm ta}\simeq 1.06$.
@@ 5 Eulerian mapping
In this Section, we introduce the Lagrangian to Eulerian mapping assuming the simplest possible approximation: the spherical map. With this assumption, each Lagrangian halo and void is considered evolving following the spherically symmetric gravitational collapse/expansion [12, 61, 62]. For both the halo and the void cases, the number of objects is assumed to be conserved from Lagrangian to Eulerian space. Considering the void case, due to mass conservation, the corresponding Eulerian radius is $R_{\rm E}=(1+\delta_{\rm v}^{\rm E})^{-1/3}R$, where the super- and sub-script E denotes Eulerian quantities. For haloes, the spherical symmetric collapse implies that all fluid elements of a Lagrangian halo collapse in the Eulerian space, therefore $M_{\rm E}(R)=M_{\rm TH}(R)$ as defined in Sec. 2. Note that the multiplicity function is a differential quantity; therefore, to express it as a function of Eulerian quantities, we use the relation
$f(S)\,{\rm d}S=f(R)\,{\rm d}R=f(R_{\rm E})\,{\rm d}R_{\rm E}=f(M_{\rm E})\,{\rm d}M_{\rm E}.$ (5.1)
Figure 4: Top panel: VSF measurements from DEMNUni (black dots) and HR-DEMNUni (red dots) at $z=0$, for voids reaching $\delta^{\rm E}_{\rm v}=-0.409$ threshold, corresponding to the linear threshold $\delta_{\rm v}=-0.623$ when the spherical symmetric evolution is assumed. Blue curves are the theoretical VSF in Eulerian space, Eq. (5.2), obtained assuming: spherical symmetry in void evolution, using the effective moving barrier Eqs. (4.1)–(4.2) for $\delta_{\rm v}=-0.623$, with the corresponding numerical multiplicity function (blue solid); Eqs. (3.4)–(3.5), and the two analytical approximations Eq. (3.14) (dashed blue) and Eq. (3.2) (dotted blue). Eulerian SVdW model (solid green) and Vdn model (solid pink). Bottom panel: relative difference between the theoretical models appearing in the upper panel (same colours) with respect to the VSF measured in DEMNUni (dark colours) and HR-DEMNUni (transparent). The errorbars show the corresponding Poissonian uncertainty.
5.1 The cosmic void case
From the above equation it follows that the Eulerian multiplicity function is $f_{\rm E}(R_{\rm E})=f[R(R_{\rm E})]\,{\rm d}R/{\rm d}R_{\rm E}=(1+\delta_{\rm v}^{\rm E})^{1/3}f[R(R_{\rm E})]$. The Eulerian void size function, therefore, is
$\frac{{\rm d}n(R_{\rm E})}{{\rm d}R_{\rm E}}=\frac{3}{4\pi R^{3}(R_{\rm E})}f_{\rm E}(R_{\rm E}).$ (5.2)